OpenStudy (jack1):

hey guys, i saw this posted earlier, actually interested on how to solve it, líl help workin it out would be appreciated please? Find the number of positive integers n less than 1000 such that 20^n - 13^n - 7^n is divisible by 309

OpenStudy (jack1):

so would u use logs, or is there a formula, or a simpler logical way?

OpenStudy (pawanyadav):

I think its not by log because we can't take log inside the brackets Log(20^n-13^n-7^n)

OpenStudy (jack1):

hmmm... ok, so what would u recommend going with? thought maybe working out the primes of 309 (3 and 103)... then using that somehow...?

OpenStudy (pawanyadav):

What is (13+7)^n I think its first and last term will cancelled 13^n and 7^n

OpenStudy (jack1):

form what i remember tho: that's not a rule of exponentials, is it?

OpenStudy (pawanyadav):

Isn't it binomial

OpenStudy (jack1):

well... for example: if n=3 13^3 +7^3 = 2540 (13+7)^3 = 20^3 = 8000

OpenStudy (pawanyadav):

No. What is binomial theorem

OpenStudy (pawanyadav):

(13+7)^3=13^3+3(13^2)(7)+3(7^213)+7^3

OpenStudy (jack1):

ah... my bad... but they're like terms(no variables)...we cant add them in this scenario?

OpenStudy (pawanyadav):

Because I want to cancel the (-13^n-7^n) part of question

OpenStudy (jack1):

i understand... but that still doesn't seem to math properly in my head... sorry :(

OpenStudy (pawanyadav):

And we know 13^n will be first and 7^n will be last term of (13+7)^n

OpenStudy (pawanyadav):

20^n- 13^n -7^n =(13+7)^n- 13^n -7^n =13^n + middle part of expansion+7^n -13^n -7^n =middle part of expansion In this way we are left with middle part of expansion We have to find any relation between this part and 309

OpenStudy (agreene):

I'm interested to see how this resolves. Because the way I would do this would be to just write a program to tell me lol

OpenStudy (pawanyadav):

Do you know. (a + b)^n = ??

OpenStudy (jack1):

yep, ur totally right, I'm with u now, sorry

OpenStudy (pawanyadav):

Its not working ,I try it with pen and paper very confused

OpenStudy (pawanyadav):

@ganeshie8 may help

OpenStudy (baru):

i think @mukushla knows how to solve this.